Report of Activities on the RVIB N.B. Palmer
21 August 2001

Thinner ice pack, clear skies with unlimited visibility, and a number of open leads made traveling easier today (22 August) as we moved to the outer margin of the continental shelf from station #87 on the inner shelf. The outer shelf region on the southern portion of the broad-scale survey grid is the last to be sampled during this cruise and then only in an abbreviated fashion, given the need to move to the Gould's location by mid-day on the 24th of August and then assist that ship move up the coast to Palmer Station (-67° 21S; -70° 26W). Our current position at 2357 is -68° 02.795S; -74° 43.787W. Winds are out of the southwest (227) at 15 to 22 kts and the air temperature is -14.2°C. The barometer reads 969.9 mlb and is slowly rising.

The pack ice had its way with us on 21 August. Our attempts at getting to two of the scheduled stations close to the Charcot Island shoreline proved to be too difficult. We had already come to the realization that station #85 was a pipe dream and had not included it in the schedule. But heavy ice ridging and thick floes made it too time consuming to push on to stations #84 or 86. Thus, part way to station #84, we opted to turn and steam instead to station #87. During this transit, high winds out of the east in the 45 to 55 kt range continued until early morning. There was one gust recorded on the bridge around 0100 that was 100 kts.

In the early morning light, during the approach to the station, we came across an area with a lot of big open leads. BIOMAPER-II was deployed in hopes of getting a short run to characterize the backscattering environmental and to get a VPR profile of zooplankton images, but an electronic failure dashed those hopes. The site looked good, however, for other towing, so in spite of some fairly tough ridges, the decision was made to try and make a run of about 4 nm through a couple of the leads and then tow the MOCNESS and Tucker Trawl systems in order to characterize the zooplankton and micronekton in this our most southern station on the survey grid. The initial attempt in late morning to start a 1-m2 MOCNESS tow was ended after an encounter with ice in the wake during the launch. But a second attempt in the early afternoon proved better, but ended with some difficulties as described below and led to another tow at the end of the station. A 10-m2 MOCNESS was deployed successfully and made the entire run with only one wire snag that required some corrective action. While temperatures on deck did not feel all that cold because winds were light during the day, the sensors for both MOCNESS systems froze up while on deck and once in the water took about 30 minutes to thaw out. "Flying" a MOCNESS without good pressure readings is disconcerting given the complexities of towing in the ice pack.

A live collection to gather animals for experimental purposes was done next with the Tucker Trawl. As a result of BIOMAPER-II being incapacitated, the two frequency HTI echosounder was deployed in the early evening from the starboard stern crane in a quiet area of a lead to get a time-series acoustic record while Plummet net casts were done to ground truth the water column. A CTD cast followed a couple of hours of ice collecting and by midnight, the Palmer was being positioned to enable an under ice survey with the ROV to be done in an area undisturbed by the ship. A final task in the late night period was to use the 1-m2 MOCNESS to take a short "repeat" tow in the upper 100 m.

John Klinck reports on 21 August, the CTD group did one cast today in the middle of the 060 line. This cast could not be done to the bottom as all acoustic transducers were obscured by ice. We did not know the depth nor could we see the bottom pinger. So the cast was done to 300 m even though the bottom might have been 400 or more meters deep.

Station #87 (? m) showed a uniform property mixed layer to 90 m. The pycnocline extended over the next 100 m. There are a number of 5 to 10 m thick constant property layers in the pycnocline. Below 200 m, the oxygen is almost constant, but the temperature and salinity were both increasing. No interior extrema were observed.

Ana Sirovic reported that she deployed only one sonobuoy per day on 20 and 21August. The first one was deployed as the Palmer was backing and ramming through the pack ice and the second one was placed in the lead that we were working in for most of the day on the 21st (station #71). Although they were deployed under very different circumstances, they were both short lived and reception was lost after about 45 minutes. In the afternoon of the 21st, Jeff Otten and Ana climbed the science mast to turn the Sinclair antenna around. It was turned to ~45° angle away from the stern, most likely due to the heavy wind that was blowing during the night. This may have been the reason for quick loss of the sonobouy deployed on the 21st. Work has also been proceeding on making wav file recordings of some seal calls recorded during the cruise.

BIOMAPER-II/MOCNESS report (P. Wiebe, C. Ashjian, and S. Gallager):
The morning of 21 August was not a particularly good one from a BIOMAPER-II sampling point of view. The sea ice conditions, with low winds and large leads in the area, were better than had been seen in quite a while and so the towed body was deployed about 0900. Normally, it is deployed with the main power turned on to keep the electronics in the telemetry bottle above the freezing mark and the other electronic sub-systems are powered up once the towed body is in the water. This time when power to the acoustic system was turned on, the rest of the system failed. When the power was turned off to the HTI system, the other sub-systems would run. The towed body was recovered about 45 minutes later, when all the testing that could be done from the surface had run its course and it was clear an electronic problem existed that required access to the electronics in the under water housings to troubleshoot. Later in the day, it was discovered that a 48-volt power supply for the acoustic system had shorted out and needed to be replaced. Although a replacement part is not available on the Palmer, a possible work-around is being explored in an attempt to get the system running again.

Two 1-m2 MOCNESS tows were conducted at station #87. A first attempt to do tow #14 did not succeed. During the launch, with the net barely in the water, the bridge, apparently thinking it was OK to kick up the pitch, sent the net sailing back with the propeller wash and driving surface ice chunks into the net. In spite of the MT instantly reacting to the problem and requesting a full stop, the bridge reaction time was not fast enough to prevent the net system from getting its share of bumps against the ice. It took a few minutes to get the net and towing wire clear of the ice and then it was brought back on board. Net zero had a big load of slush in it and the flow meter was not functional. So with the net raised up off the deck, the ice was moved forward in the net and dumped out onto the deck in front of the net frame. The flow meter suffered some damage and was replaced. The net bar traps also needed re-adjustment. Fortunately, there was no other damage.

The second attempt at tow #14 was done at approximately noon local time. The ship was positioned in a long lead (~4 miles) fashioned by breaking ice between two smaller leads. The tow was conducted to 309 meters, at a bottom depth of approximately 360 m. During the upcast portion of the tow, the lead started to close in and it became difficult for the ship to maintain a steady speed. Ice chunks were cast off and into the wake of the ship, resulting in the wire being caught on several large chunks, which resulted in the ship stopping to free the wire. One particularly persistent chunk resulted in the MOCNESS ascending rapidly from approximately 50 m to 10 m, or just below the ice pack undersurface. The MOCNESS pressure sensor apparently froze during the tow and was giving incorrect readings at this time, however we were able to use the pressure from the OPC, which was mounted on the MOCNESS to determine the depth. After this near-ice incident, the tow was aborted and the net brought on board fairly quickly. As a result, the last two nets (50-25 m and 25-0 m depth intervals) did not fish. The pressure sensor read 20 m when the net was at the surface. Also, the angle was set incorrectly and was 9° greater than it should have been during the tow; the volumes of the nets will have to be recalculated using the corrected angles to obtain a realistic net mouth area.

The second tow (Tow #15) was conducted as the last activity at this very long station and was done to sample the upper water column where it had been impossible to sample during Tow 14. For this tow, the ship ran 1.5 miles along a lead, then back to the starting point, in order to see if there was sufficient open water to conduct the tow. The tow was conducted at approximately 0330 on 22 August and sampled two depth intervals: 100-50 m and 50-5 m. We did not run out of open water during the tow and were able to deploy and recover the net in the open water of the lead. Because of the low ship speed (1.5 knots) necessary to complete sampling within open water, the net angle was somewhat lower than optimal, being in the 20-30° range. The pressure sensor did not function upon deployment, so the OPC depth sensor was used to achieve a net depth of 100 m and to determine when to trip the first net. However, the MOCNESS pressure sensor started to work shortly thereafter and continued to function for the remainder of the tow.

Copepods were dominant throughout the water column and were especially abundant at the lower depths, from 200-300 m. They were substantially more abundant at this location than at all previous sampling locations. Krill furcilia were collected from 150-200 m. Siphonophores were seen at the bottom of the water column (250-309 m). Thysanoessa larvae were observed over most of the water column. Ctenophores were observed n the upper 50-75 m, and were very abundant in the upper 50 m.

Cheers, Peter